OpenStudy (anonymous):

Multiply and simplify: (2a – 1)(2a + 1) A. 4a2 – 1 B. 4a2 + 1 C. 4a2 + -4a – 1 D. 4a2 + 4a – 1

Parth (parthkohli):

You can do this in two ways. * (a + b)(a - b) = a^2 - b^2 * Really multiply them

OpenStudy (anonymous):

Use this formula: \[(a-b)(a+b) = a^2 - b^2\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2a(2a + 1) - 1(2a + 1)}\) \( \color{Black}{\Rightarrow 4a^2 + 2a - 2a - 1}\)

OpenStudy (anonymous):

B ?

OpenStudy (theviper):

4a^2+2a-2a-1

Parth (parthkohli):

Na, it's -1 not +1

OpenStudy (anonymous):

Check it again It is not B...

OpenStudy (theviper):

It's A.

Parth (parthkohli):

I prefer really multiplying them for understanding purposes

OpenStudy (anonymous):

Thanks viper for making it easy for the senior...

OpenStudy (theviper):

yw:)

OpenStudy (theviper):

LOL haha

OpenStudy (anonymous):

@TheViper : BECOME MY FAN SO I CAN MESSAGE YOU.

OpenStudy (theviper):

k